OpenStudy (anonymous):

Find the equation of the line that runs parallel to x-5=0 and passes through the point (8,6).

OpenStudy (anonymous):

how do i find slope?

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

no slope, line is vertical

OpenStudy (anonymous):

what tracy said

OpenStudy (anonymous):

so how would i write the equation?

OpenStudy (anonymous):

just have no slope?

OpenStudy (anonymous):

the slope is 0 because x-5=0 is a vertical line. You would write the parallel line as x=8

OpenStudy (anonymous):

but if it asks for the equation just do y-6=(x-8) or is there no equation because there is no slope

OpenStudy (anonymous):

the equation y-6=(x-8) has a slope of 1

OpenStudy (anonymous):

@tracy slope of course is not zero it is undefined

OpenStudy (anonymous):

slope of \[y=2\] is zero slope of \[x=5\] does not exist

OpenStudy (anonymous):

oh yea sorry, horizontal lines have slope=0

OpenStudy (anonymous):

so that line is undefined because its slope is non existent?

OpenStudy (anonymous):

there is no equation of that line?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x-5%3D0

OpenStudy (anonymous):

It's not a vertical line lol...my bad

OpenStudy (anonymous):

no it has an equation. the equation for vertical line looks like \[x=\text{some number}\]

OpenStudy (anonymous):

it just doesn't have a slope, so you cannot for example use the point slope formula

OpenStudy (anonymous):

so the equation would simply be x=8

OpenStudy (anonymous):

no, it's y = x-5, with a slope of 1/1x - 5, -5 is the y-intercept.

OpenStudy (anonymous):

y = mx+b m is slope, b is y-intercept. we can put this equation in: 0 = 1x+(-5) y = 0, m = 1, b = -5.

OpenStudy (anonymous):

No, because y is a variable. The original equation has x=5 even when y is not 0. Your equation works at that point, but is not parallel to x-5 = 0.